a merchant has coffee worth $10 a pound that she wants to mix with 40 pounds of coffee worth $80 a pound to get a mixture that can be sold for $30 a pound how many pounds of $10 coffee should be used?
Answer by checkley79(3341)

10x+80*40=30(x+80)
10x+3200=30x+2400
10-30x=2400-3200
-20x=-800
x=-800/-20
x=40 lbs. of $10 coffee is used.
Proof:
10*40+80*40=30(40+80)
400+3200=30*120
3600=3600